M Chapter 17 Hw 17-4jpg ntpx 10.00 points Problem 17-19 Preferred stock dividends in arrears (L017-51 Robbins Petroleum Company is four years in arrears on cumulative preferred stock dividends. There are 880,000 prefermed shares outstanding, and he aneual dvidend is $9 50pe share. The Vice-President of Finance sees nor a hope of paying the di den s n a rears. She S devising a plan to compe sa ete pretened stock ders tr poi en t dividends in arrears. a. How much should the compensation be? (Do not round intermediate calculations, Input your answer in dollars, not millions (o.g. $1,.234,000 bonds. The bonds will have a 15-year maturity. Using the bond valuation Table 16.2. indicate the market value of a $1,000 par value bond (Round your answer to the nearest whe number.) t in which the going rate ot interest is 16 percent for sim O Type here to search
M Chapter 17 HW tpx umber c. Based on market value, how many bonds must be issued to provide the compensation determined in part a? (Do not round intermediate calculations and round y Worksheet stock dividends in arrears LO17-5 stock is an that falls somewhere between debt and type of
Answer #1

a.Annual Dividend per share = $9.50

Number of shares = 880,000

Dividend Outstanding = 880,000*9.50*4 = $33,440,000

Compensation = 90%

i.e. = $33,440,000*90%

= $30,096,000

b.Market value of bond is equal to the present value of all interest and principal payments at market rate of interest

Market Value = 120*PVAF(18%, 15 years) + 1000*PVF(18%, 15 years)

= 120*5.092 + 1,000*0.084

= $695.04

c.Number of bonds required to be issued = $30,096,000/695.04

= 43,301 (approx.)